# I Know What I Don’t Know: Improving Model Cascades Through Confidence Tuning

### v1

@@ Abstract
Large-scale machine learning models deliver strong performance across a wide range of tasks but come with significant computational and resource constraints. To mitigate these challenges, local smaller models are often deployed alongside larger models, relying on routing and deferral mechanisms to offload complex tasks. However, existing approaches inadequately balance the capabilities of these models, often resulting in unnecessary deferrals or sub-optimal resource usage. In this work we introduce a novel loss function called Gatekeeper for calibrating smaller models in cascade setups. Our approach fine-tunes the smaller model to confidently handle tasks it can perform correctly while deferring complex tasks to the larger model. Moreover, it incorporates a mechanism for managing the trade-off between model performance and deferral accuracy, and is broadly applicable across various tasks and domains without any architectural changes. We evaluated our method on encoder-only, decoder-only, and encoder-decoder architectures. Experiments across image classification, language modeling, and vision-language tasks show that our approach substantially improves deferral performance.
@@ 1   Introduction
Figure 1: Overview of the cascading setup (left) and performance trade-off (right). Left: Cascading determines which inputs should be predicted by a small model $\mathcal{M}_{S}$ or routed to a large model $\mathcal{M}_{L}$. Right: Performance is measured as a trade-off between joint accuracy across $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$ and deferral ratio. Ideal deferral strategies optimize this trade-off and push the realized deferral curve closer to the ideal deferral depicted in (d). (a) depicts full deferral; (b) depicts no deferral; and (c) depicts excessive deferral of requests that could have been correctly handled by $\mathcal{M}_{S}$.
@@
In recent years, large-scale machine learning models such as Gemini (GeminiTeam et al., 2023), GPT-4 (Achiam et al., 2023) or Claude (Anthropic, 2024) have gained significant traction due to their remarkable ability to address a wide array of tasks. These tasks range from natural language understanding and generation, including machine translation, summarization, and conversational agents, to computer vision applications like image recognition, object detection, and image captioning. The versatility and high performance of these expansive models make them invaluable tools across diverse domains, including healthcare (Nazi & Peng, 2024), finance (Li et al., 2023), education (Wang et al., 2024b), and entertainment (Gallotta et al., 2024).
Deploying and operating such large models presents significant challenges in terms of latency, memory, compute and storage (Pope et al., 2023). Optimizing inference costs is an active research area which includes both techniques for reducing the size of the existing large model such as model compression  (Hoefler et al., 2021), model pruning (Ma et al., 2023; Cheng et al., 2024) and distillation (Yang et al., 2024), and those aiming to leverage a sequence of models such as speculative decoding (Leviathan et al., 2023) and model cascades (Dohan et al., 2022; Chen et al., 2023; Gupta et al., 2024; Chen et al., 2024a). However, due to scaling laws showing that the performance of a Large Language Model (LLM) increases with its size (Kaplan et al., 2020), the latter category of methods leveraging a sequence of models is currently a more promising direction to lower inference costs without sacrificing the capabilities of large models.
Both speculative decoding and model cascading rely on the existence of a large performant model $\mathcal{M}_{L}$ and a small model $\mathcal{M}_{S}$ that is cheap, fast, and less accurate. Speculative decoding leverages $\mathcal{M}_{S}$ for generating a set of draft tokens that are then validated by $\mathcal{M}_{L}$ in parallel, a technique successfully deployed in industry applications (Leviathan, 2024). In contrast, model cascades leverage a deferral rule for selecting the most suitable model to process a given request (see Figure 1 left). While the success of the speculative decoding necessitates a highly performant $\mathcal{M}_{S}$ to generate quality draft tokens, model cascades allow the deployment of a less capable $\mathcal{M}_{S}$ by invoking $\mathcal{M}_{L}$ only for inference requests outside the small model’s scope. In this work, we contribute to the advancement of the model cascades.
Model cascades achieve efficient deferral by optimizing for two objectives: compute budget and joint accuracy. We illustrate their trade-off on the example shown in Figure 1 (right). Assume we have x inference requests and a small model $\mathcal{M}_{S}$ that only requires 20% of the compute budget of the large model $\mathcal{M}_{L}$. There are three worst case scenarios: (a) the small model $\mathcal{M}_{S}$ defers all requests to $\mathcal{M}_{L}$ and the system achieves the best joint accuracy (equal to the accuracy of $\mathcal{M}_{L}$) but the worst compute budget (1.2x) since all requests are run on both models; (b) $\mathcal{M}_{S}$ never sends a request to $\mathcal{M}_{L}$, resulting in the smallest compute budget (0.2x) but also the lowest joint accuracy (equal to the accuracy of $\mathcal{M}_{S}$); (c) $\mathcal{M}_{S}$ only sends requests that it could have answered correctly to $\mathcal{M}_{L}$, requiring an increased compute budget compared to (b) but still resulting in the lowest joint accuracy (equal to the accuracy of $\mathcal{M}_{S}$). On the other hand, an ideal case is the scenario (d) where the small model $\mathcal{M}_{S}$ only sends requests for which it would be incorrect, requiring a compute budget between 0.2-1x but resulting in the optimal joint accuracy given that budget. We call the approximation of the ideal case the deferral performance.
In this paper we address the following research question:
How can we optimize model cascades to maximize deferral performance?
In other words, we focus on designing effective model cascades by making the small model more aware of what it does not know. We achieve this by introducing a general-purpose loss function, called Gatekeeper, that calibrates the small model’s confidence in its predictions. By fine-tuning $\mathcal{M}_{S}$ to output high confidence for correct predictions and low confidence for incorrect ones, we enhance the reliability of its uncertainty estimates and facilitate learning of common tasks, thereby directly improving the deferral performance. Crucially, Gatekeeper includes an inherent mechanism for managing the trade-off between model performance and deferral performance that can be applied to an arbitrary architecture, making our work directly applicable to Vision-Language Models (VLMs).
We demonstrate the efficacy of the Gatekeeper loss across various model architectures, including encoder-only vision models for image classification, decoder-only LMs for closed- and open-form text generation, and encoder-decoder setups for VL tasks for open set classification and captioning. Our main results show that models trained with Gatekeeper outperform an untuned baseline by a factor of 0.72x/2x on CIFAR-100/TinyImagenet and 7x/10x on ARC-e/c, respectively, in terms of deferral performance. This advancement paves the way for more scalable and efficient deployment strategies, leveraging the strengths of both local and large-scale models to deliver high-quality results in real-time applications.
@@ 2   Related Work
Our proposed method improves model cascades through uncertainty-aware finetuning. Next, we describe related work for both research areas.
@@
Model Cascades: A cascade consists of a series of models and a deferral rule which determines the appropriate model given an input request. The concept of model cascades has first been proposed by Viola & Jones (2001), where it is used to accelerate object detection models. Cascades have been extensively studied for classification-based computer vision (Wang et al., 2017; Trapeznikov & Saligrama, 2013; Bolukbasi et al., 2017; Jitkrittum et al., 2023) and in models for natural language processing (Dohan et al., 2022; Mamou et al., 2022; Varshney & Baral, 2022).
Cascades are particularly promising in the context of generative models such as LLMs and VLMs since they can significantly reduce inference costs. In contrast to speculative decoding (Leviathan et al., 2023), they aim to invoke the large model only for difficult examples. However, the two approached can also be combined (Narasimhan et al., 2025). While Chen et al. (2024b) combine the deferral logic with speculative decoding to generate initial tokens using larger models and later tokens using a smaller model, the majority of research on model cascades has focused on using pre-trained LLMs with a post-hoc deferral logic (Narasimhan et al., 2022; Jitkrittum et al., 2023; Yue et al., 2024). Kolawole et al. (2024) use agreement across multiple models to make deferral decisions, while Gupta et al. (2024) present a method to learn a deferral rule based on quantiles of per-token log probabilities.
Model cascades can be further improved through training and fine-tuning. Wang et al. (2024a) train the small model only on easier examples by masking tokens for which large and small model are incorrect. Enomoro & Eda (2021) extend the training objective of image classification models with confidence calibration. In contrast to previous research, our work extends cascades to VLMs and improves overall inference performance by making smaller models less confident when they are incorrect.
Uncertainty-Aware Models: Extensive research has been conducted in the field of uncertainty quantification in deep learning and we refer to Abdar et al. (2021) for a detailed survey. While many methods have been proposed for classification-based models, measuring uncertainty for generative models is still an active area of research. Based on the assumed level of access to model internals, existing methods can be summarized into three main categories:
Black box methods operate solely via the model’s query interface by injecting tailored instructions into prompts. These modify the prompt $\mathbf{x}$ by appending instructions $\mathbf{x}^{\prime}$ for the model to respond less confidently: $\mathbf{x}\leftarrow\mathbf{x}|\mathbf{x}^{\prime}$. Related methods are confidence quantification (Shrivastava et al., 2023), rejection and remote model awareness (Kadavath et al., 2022), and self-critiquing (Gou et al., 2023). Xiong et al. (2024) show that LLMs can express their confidence through prompting and sampling strategies and their experiments indicate that these models tend to be overconfident.
Gray box approaches employ confidence-based strategies centered on post-processing the model’s logits. Many uncertainty techniques such as ensembling (Lakshminarayanan et al., 2017) and Bayesian methods (Blundell et al., 2015)) are not scalable. Related techniques are max confidence (Hendrycks & Gimpel, 2016), predictive entropy, and confidence reduction prompting. Malinin & Gales (2021) uses token-entropy as a measure of uncertainty in auto-regressive models and Kuhn et al. (2023) leverages linguistic invariances via semantic entropy.
White box methods utilize uncertainty-aware fine-tuning in order to produce more accurately calibrated models. Chuang et al. (2024) introduces Self-REF, a framework which leverages confidence tokens during fine-tuning to improve performance in downstream routing. Krishnan et al. (2024) proposes an uncertainty-aware causal language modeling loss function, which captures the trade-off between predictive accuracy and uncertainty calibration. In contrast to previous work, our method aims to calibrate the model in a way such that correctly generated predictions are assigned low predictive uncertainty and incorrectly generated predictions are assigned high predictive uncertainty. We consider the uncertainty-aware model in the context of cascade inference system, where it helps to improve overall performance.
Prior to this work, Rawat et al. (2021) has proposed pre-partitioning the data into easy and hard sets before training. One example of such a partitioning is based on the $\mathcal{M}_{L}$’s confidence. The small model $\mathcal{M}_{S}$ is then trained to identify easy/hard examples with explicit labels. We improve on this static one-time partitioning of easy/hard examples by deciding the assignment dynamically based on $\mathcal{M}_{S}$’s current state during training.
3   Method
@@ 3.1 Overview & Setup
Our framework consists of a large, highly capable model $\mathcal{M}_{L}$ and a smaller, resource-efficient model $\mathcal{M}_{S}$. We assume that $S\in\mathbb{N}$ and $L\in\mathbb{N}$ represent the parameter count of each model with $S\ll L$. Both models can either function as classifiers (i.e., $\mathcal{M}:\mathbb{R}^{D}\rightarrow[C]$ with $\mathbb{R}^{D}$ denoting the input space and $C$ the number of total classes), or (multi-modal) sequence models (i.e., $\mathcal{M}:\mathbb{R}^{D}\rightarrow[V]^{T}$ where $V$ is the vocabulary and $T$ is the sequence length). We include experiments on all of these model classes in Section 4. Furthermore, we do not require a shared model family to be deployed on both $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$; for example, $\mathcal{M}_{S}$ could be a custom convolutional neural network optimized for efficient inference and $\mathcal{M}_{L}$ a vision transformer (Dosovitskiy, 2020). The primary objective is to design a deferral mechanism that enables $\mathcal{M}_{S}$ to decide when to return its predictions without the assistance of $\mathcal{M}_{L}$ and when to instead defer to it.
Deferral decisions are made using signals derived from the small model $\mathcal{M}_{S}$ as this approach is typically more cost-effective than employing a separate routing mechanism (Teerapittayanon et al., 2016). Approaches that involve querying the large model $\mathcal{M}_{L}$ to assist in making deferral decisions at test time are excluded from our setup. Such methods — common in domains like LLMs — are counterproductive to our goal since querying $\mathcal{M}_{L}$ defeats the purpose of making a deferral decision in the first place?. Examples of these inapplicable methods include collaborative LLM frameworks (Mielke et al., 2022) and techniques that rely on semantic entropy for uncertainty estimation (Kuhn et al., 2023). As part of our setup, we assume that $\mathcal{M}_{S}$ is strictly less capable than $\mathcal{M}_{L}$— a realistic scenario in practice supported by scaling laws (Kaplan et al., 2020). Under this assumption, mistakes made by $\mathcal{M}_{L}$ are also made by $\mathcal{M}_{S}$; however, $\mathcal{M}_{S}$ may make additional errors that $\mathcal{M}_{L}$ would avoid. This reflects the general observation that larger models tend to outperform smaller models across a wide range of tasks.
As discussed in Section 2, the choice of deferral strategy often depends on the level of access available to $\mathcal{M}_{S}$. We assume white box access with full access to $\mathcal{M}_{S}$’s internals. As such, deferral mechanisms can be directly integrated into the model’s architecture and parameters. This involves fine-tuning $\mathcal{M}_{S}$ to predict deferral decisions or to incorporate rejection mechanisms within its predictive process. Our work falls into this category as it proposes a new loss function to fine-tune $\mathcal{M}_{S}$.
Our goal is to train a small model that can effectively distinguish between correct and incorrect predictions. While many past works have considered the question of whether it is possible to find proxy measures for prediction correctness, the central question we ask is:
Can we optimize the small model $\mathcal{M}_{S}$ to separate correct from incorrect predictions?
We show that this is indeed achievable through a carefully designed fine-tuning stage that does not require any architectural modifications. This ensures that the ability to separate correct from incorrect decisions is integrated seamlessly into $\mathcal{M}_{S}$’s existing structure.
@@ 3.2 Confidence-Tuning for Deferral
CorrectIncorrectBatchClass 1Class 2Class 3Class 4Class 5000.50.50.50.51111ProbabilityMinimize $\text{CE}({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}p_{i}},y_{i})$Class 1Class 2Class 3Class 4Class 5000.50.50.50.51111ProbabilityMinimize $\text{KL}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{i}}\parallel\mathcal{U}\right)$
Figure 2: Overview of Gatekeeper: We want correctly predicted samples maintain their current prediction by ensuring that cross entropy is decreased (top, green). At the same time, we want incorrectly predicted samples to yield a uniform confidence across all classes, leading to a low overall confidence score (bottom, red).
Stage 1: Standard Training. We begin with a $\mathcal{M}_{S}$ that has already been trained on the tasks it is intended to perform upon deployment. However, due to its limited capacity, $\mathcal{M}_{S}$ cannot achieve the performance levels of $\mathcal{M}_{L}$. Importantly, we make no assumptions about the training process of $\mathcal{M}_{S}$ —whether it was trained from scratch without supervision from an external model or through a distillation approach.
Stage 2: Correctness-Aware Finetuning with Gatekeeper. Next, we introduce a correctness-aware loss, dubbed Gatekeeper, to fine-tune $\mathcal{M}_{S}$ for improved confidence calibration. Specifically, the model is trained to make correct predictions with high confidence while reducing the confidence of incorrect predictions (see Figure 2). This loss can either rely on true labels or utilize the outputs of $\mathcal{M}_{L}$ with soft probabilities as targets.
For a standard classification model, the calibration loss is defined as the following hybrid loss
@@
$\mathcal{L}_{\text{corr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\{y_{i}=\hat{y}_{i}\}\text{CE}(p_{i}(\mathbf{x}_{i}),y_{i})$ (2)
$\mathcal{L}_{\text{incorr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\{y_{i}\neq\hat{y}_{i}\}\text{KL}\left(p_{i}(\mathbf{x}_{i})\parallel\mathcal{U}\right)$ (3)
where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted labels for $\mathbf{x}_{i}$, respectively, $p_{i}$ is the predicted probability distribution of $\mathcal{M}_{S}$ over classes, $\mathcal{U}$ represents the uniform distribution over all classes, $N$ denotes the number samples in the current batch, $\alpha\in(0,1)$ is a tunable hyperparameter controlling the emphasis between correct and incorrect predictions, and the cross-entropy function and KL divergence are defined as $\text{CE}(p,y)=-\sum_{c}y_{c}\log p_{c}$ and $\text{KL}(p\parallel q)=\sum_{c}p_{c}\log(\frac{p_{c}}{q_{c}})$, respectively. We note that a similar loss has previously been proposed in Outlier Exposure (OE) (Hendrycks et al., 2018) for out-of-distribution (OOD) sample detection. Here, the goal is to make sure that OOD examples are assigned low confidence scores by tuning the confidence on a auxiliary outlier dataset. However, to the best of our knowledge, this idea has not previously been used to improve deferral performance of a smaller model in a cascading chain.
We emphasize that the trade-off parameter $\alpha$ plays a critical role as part of this optimization setup as it directly influences model utility and deferral performance. A lower value of $\alpha$ emphasizes reducing confidence in incorrect predictions by pushing them closer to the uniform distribution, making the model more cautious in regions where it may make mistakes. Conversely, a higher value of $\alpha$ encourages the model to increase its confidence on correct predictions, sharpening its decision boundaries and enhancing accuracy where it is already performing well. Thus, $\alpha$ serves as a crucial hyperparameter that balances the trade-off between improving calibration by mitigating overconfidence in errors and reinforcing confidence in accurate classifications. By appropriately tuning $\alpha$, practitioners can control the model’s behavior to achieve a desired balance between reliability in uncertain regions and decisiveness in confident predictions, tailored to the specific requirements of their application.
We further generalize this loss to token-based models (e.g., LMs and VLMs), formulated as
$\mathcal{L}_{\text{corr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathds{1}\{y_{i,t}=\hat{y}_{i,t}\}\text{CE}(p_{i,t}(\mathbf{x}_{i}),y_{i,t})$ (4)
$\mathcal{L}_{\text{incorr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathds{1}\{y_{i,t}\neq\hat{y}_{i,t}\}\text{KL}\left(p_{i,t}(\mathbf{x}_{i})\parallel\mathcal{U}\right)$ (5)
where $y_{i,t}$ and $\hat{y}_{i,t}$ denote the true and predicted tokens at position $t$ for sample $i$, $p_{i,t}$ is the predicted token distribution at position $t$ for sample $i$, and $T$ is the sequence length for the token-based model. The token-level loss ensures that correct token predictions are made confidently while incorrect tokens are assigned smaller confidences.
Stage 3: Confidence Computation & Thresholding. After fine-tuning $\mathcal{M}_{S}$ with Gatekeeper, we apply standard confidence- and entropy-based techniques for model uncertainty to obtain a deferral signal. We use the selective prediction framework to determine whether a query point $\mathbf{x}\in\mathbb{R}^{D}$ should be accepted by $\mathcal{M}_{S}$ or routed to $\mathcal{M}_{L}$. Selective prediction alters the model inference stage by introducing a deferral state through a gating mechanism (El-Yaniv & Wiener, 2010). At its core, this mechanism relies on a deferral function $g:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which determines if $\mathcal{M}_{S}$ should output a prediction for a sample $\mathbf{x}$ or defer to $\mathcal{M}_{L}$. Given a targeted acceptance threshold $\tau$, the resulting predictive model can be summarized as:
$(\mathcal{M}_{S},\mathcal{M}_{L},g)(\mathbf{x})=\begin{cases}\mathcal{M}_{S}(\mathbf{x})&g(\mathbf{x})\geq\tau\\ \mathcal{M}_{L}(\mathbf{x})&\text{otherwise.}\end{cases}$ (6)
Classification Models (Max Softmax). Let $\mathcal{M}_{S}$ produce a categorical distribution $\{p(y=c\mid\mathbf{x})\}_{c=1}^{C}$ over $C$ classes. Then we define the gating function as
$g_{\text{CL}}(\mathbf{x})\;=\;\max_{1\,\leq\,c\,\leq\,C}\;p\bigl{(}y=c\,\big{|}\,\mathbf{x}\bigr{)}.$ (7)
Token-based Models (Negative Predictive Entropy). Let $\mathcal{M}_{S}$ produce a sequence of categorical distributions $\{p(y_{t}=c\mid\mathbf{x})\}_{c=1}^{C}$ for each token index $t\in T$. Then we define the gating function as
$\footnotesize g_{\text{NENT}}(\mathbf{x})=\;\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{C}p\bigl{(}y_{t}=c\,\big{|}\,\mathbf{x}\bigr{)}\,\log p\bigl{(}y_{t}=c\,\big{|}\,\mathbf{x}\bigr{)},$ (8)
where $y_{t}\in[C]$ is the predicted token at time step $t$, $p(y_{t}=c\mid\mathbf{x})$ is the (conditional) probability of token $k$ at step $t$, and $T$ is the total number of token positions for the sequence. Across both model classes, higher values of either $g_{\text{CL}}$ or $g_{\text{NENT}}$ indicate higher confidence in the predicted class or sequence generation, respectively.

### v2

@@ Abstract
Large-scale machine learning models deliver strong performance across a wide range of tasks but come with significant computational and resource constraints. To mitigate these challenges, local smaller models are often deployed alongside larger models, relying on routing and deferral mechanisms to offload complex tasks. However, existing approaches inadequately balance the capabilities of these models, often resulting in unnecessary deferrals or sub-optimal resource usage. In this work we introduce a novel loss function called Gatekeeper for calibrating smaller models in cascade setups. Our approach fine-tunes the smaller model to confidently handle tasks it can perform correctly while deferring complex tasks to the larger model. Moreover, it incorporates a mechanism for managing the trade-off between model performance and deferral accuracy, and is broadly applicable across various tasks and domains without any architectural changes. We evaluate our method on encoder-only, decoder-only, and encoder-decoder architectures. Experiments across image classification, language modeling, and vision-language tasks show that our approach substantially improves deferral performance.
@@ 1 Introduction
Figure 1: Overview of the cascading setup (left) and performance trade-off (right). Left: Cascading determines which inputs should be predicted by a small model $\mathcal{M}_{S}$ or routed to a large model $\mathcal{M}_{L}$. Right: Performance is measured as a trade-off between joint accuracy across $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$ and deferral ratio. Ideal deferral strategies optimize this trade-off and push the realized deferral curve closer to the ideal deferral depicted in (d). (a) depicts full deferral; (b) depicts no deferral; and (c) depicts excessive deferral of requests that could have been handled by $\mathcal{M}_{S}$.
@@
Large-scale machine learning models such as Gemini (GeminiTeam et al., 2023), GPT-4 (Achiam et al., 2023), and Claude (Anthropic, 2024) have demonstrated remarkable capabilities across diverse tasks, including language understanding, generation, and computer vision. Their strong generalization has enabled deployment in domains such as healthcare (Nazi & Peng, 2024), finance (Li et al., 2023), education (Wang et al., 2024b), and entertainment (Gallotta et al., 2024). However, the substantial computational, memory, and latency costs associated with these models pose significant scalability issues (Pope et al., 2023).
To address these challenges, two main strategies have emerged. The first compresses the large model itself—via pruning (Ma et al., 2023), distillation (Yang et al., 2024), or sparsity techniques (Hoefler et al., 2021). The second—arguably more promising given empirical scaling laws (Kaplan et al., 2020)—maintains the large model $\mathcal{M}_{L}$ and reduces cost by selectively offloading easy inputs to a smaller, cheaper model $\mathcal{M}_{S}$. This idea underlies speculative decoding (Leviathan et al., 2023) and model cascades (Dohan et al., 2022; Gupta et al., 2024; Chen et al., 2024a), which pair a fast $\mathcal{M}_{S}$ with a more powerful $\mathcal{M}_{L}$.
In speculative decoding, $\mathcal{M}_{S}$ drafts candidate outputs that $\mathcal{M}_{L}$ verifies in parallel. In contrast, model cascades rely on a deferral rule to decide which model should handle a given request (Figure 1, left). Cascades are attractive because they permit the use of a less capable $\mathcal{M}_{S}$—provided it can accurately identify when it is likely to make mistakes. The success of such systems hinges on balancing compute cost and joint accuracy. As shown in Figure 1 (right), cascades can fail if $\mathcal{M}_{S}$ defers too often, too little, or defers only on correct predictions. The optimal scenario occurs when $\mathcal{M}_{S}$ defers only when it is wrong, yielding the best accuracy–efficiency trade-off. We refer to how well a cascade approximates this ideal as its deferral performance.
Our Contribution. We address the question: How can we optimize model cascades to maximize deferral performance? That is, how can we train $\mathcal{M}_{S}$ not only to be competent on easy examples, but also to know when to defer? We propose Gatekeeper, a simple and general-purpose loss function that fine-tunes $\mathcal{M}_{S}$ to output high confidence when correct and low confidence when incorrect. This calibration improves uncertainty estimates, enabling more accurate routing and better cascade-level performance. Our approach directly shapes $\mathcal{M}_{S}$ ’s confidence through uncertainty-aware fine-tuning. Gatekeeper also includes a tunable parameter that controls the trade-off between predictive and deferral performance, and is applicable to a wide range of architectures without architectural modifications.
We demonstrate the effectiveness of Gatekeeper on encoder-only vision models, decoder-only language models, and encoder-decoder vision-language models. Across tasks such as image classification, closed-form text generation, and image captioning, Gatekeeper significantly improves deferral performance. For instance, it achieves up to 2×2$\times$2 × improvement on TinyImageNet and 10×10$\times$10 × on ARC-e/c. These results show that Gatekeeper enables more reliable, cost-efficient cascaded inference—paving the way for scalable deployment of machine learning systems across domains.
@@ 2 Related Work
Model Cascades: Model cascades consist of a deferral rule and a sequence of models, routing inputs to the appropriate model based on difficulty. Originally introduced to accelerate object detection (Viola & Jones, 2001), cascades have since been applied in classification (Wang et al., 2017; Trapeznikov & Saligrama, 2013; Bolukbasi et al., 2017; Jitkrittum et al., 2023) and NLP (Dohan et al., 2022; Mamou et al., 2022; Varshney & Baral, 2022).
Cascades are especially useful for LLMs and VLMs, reducing inference cost by deferring only hard examples to larger models. Unlike speculative decoding (Leviathan et al., 2023), which accelerates generation, cascades focus on selective model invocation, though both can be combined (Narasimhan et al., 2025; Chen et al., 2024b). Most prior work applies post-hoc deferral logic to pre-trained models (Narasimhan et al., 2022; Yue et al., 2024; Kolawole et al., 2024; Gupta et al., 2024). Recent approaches improve deferral through training: Wang et al. (2024a) restrict $\mathcal{M}_{S}$ training to easier tokens; Enomoro & Eda (2021) introduce calibration-aware training. Our method extends these ideas to VLMs and generative models by encouraging $\mathcal{M}_{S}$ to be uncertain when wrong.
Uncertainty-Aware Models: Uncertainty estimation is well-studied for classifiers (Abdar et al., 2021), but remains challenging for generative models. Methods differ by access to model internals:
  1. 1.
    Black-box methods modify prompts to elicit more cautious responses (Shrivastava et al., 2023; Kadavath et al., 2022; Gou et al., 2023; Xiong et al., 2024).
  2. 2.
    Gray-box approaches analyze model outputs using entropy or logit post-processing (Hendrycks & Gimpel, 2016; Malinin & Gales, 2021; Kuhn et al., 2023). While effective, techniques like ensembling (Lakshminarayanan et al., 2017) and Bayesian inference (Blundell et al., 2015) are often impractical at scale.
  3. 3.
    White-box methods incorporate uncertainty into training objectives (Chuang et al., 2024; Krishnan et al., 2024). Rawat et al. (2021) pre-partition data based on $\mathcal{M}_{L}$’s confidence to train $\mathcal{M}_{S}$. In contrast, we dynamically adjust training based on $\mathcal{M}_{S}$’s uncertainty, ensuring it is confident when correct and uncertain when wrong. This improves performance in cascade setups.
3 The Gatekeeper Loss
@@ 3.1 Overview & Setup
We consider a model cascade consisting of a large, accurate model $\mathcal{M}_{L}$ and a smaller, resource-efficient model $\mathcal{M}_{S}$, with parameter counts $L\gg S$. Both models may be classifiers ($\mathcal{M}:\mathbb{R}^{D}\to[C]$) or sequence models ($\mathcal{M}:\mathbb{R}^{D}\to[V]^{T}$), and need not share the same architecture family. For instance, $\mathcal{M}_{S}$ may be a lightweight CNN and $\mathcal{M}_{L}$ a vision transformer (Dosovitskiy, 2020). Our goal is to fine-tune $\mathcal{M}_{S}$ such that it knows when to trust its predictions and when to defer to $\mathcal{M}_{L}$. Deferral signals are derived solely from $\mathcal{M}_{S}$. This avoids the additional compute cost of methods that query $\mathcal{M}_{L}$ at inference time (Mielke et al., 2022; Kuhn et al., 2023), which defeats the purpose of deferral. Our setup assumes $\mathcal{M}_{S}$ is strictly less capable than $\mathcal{M}_{L}$ —consistent with scaling laws (Kaplan et al., 2020)—so errors made by $\mathcal{M}_{L}$ are likely also made by $\mathcal{M}_{S}$, but not vice versa.
We assume white-box access to $\mathcal{M}_{S}$, enabling us to integrate the deferral mechanism into its training objective. This allows $\mathcal{M}_{S}$ to learn during fine-tuning how to distinguish between reliable and unreliable predictions. Rather than relying on post-hoc uncertainty thresholds or hand-crafted heuristics, we ask: Can we directly optimize $\mathcal{M}_{S}$ to separate correct from incorrect predictions via fine-tuning? We show this is achievable using a new loss that requires no architectural changes and integrates into standard pipelines.
@@ 3.2 Confidence-Tuning for Deferral
CorrectIncorrectBatchC1C2C3C4C5000.50.50.50.51111ProbabilityMinimize $\text{CE}({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}p_{i}},y_{i})$C1C2C3C4C5000.50.50.50.51111ProbabilityMinimize $\text{KL}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{i}}\parallel\mathcal{U}\right)$
Figure 2: Gatekeeper Overview: Correct predictions are made more confident (lower cross-entropy), while incorrect predictions are trained to produce uniform output distributions (higher entropy). This enables better deferral decisions at test time.
Stage 1: Initial Training. We begin with a $\mathcal{M}_{S}$ that has been trained on the target task. We make no assumptions about the training pipeline—$\mathcal{M}_{S}$ may be trained from scratch or via distillation.
Stage 2: Correctness-Aware Fine-Tuning. We then fine-tune $\mathcal{M}_{S}$ using a novel hybrid loss function, Gatekeeper, which promotes high confidence on correct predictions and low confidence on incorrect ones (see Figure 2). This objective is based on the intuition that a reliable deferral mechanism requires well-calibrated uncertainty estimates.
The loss takes the form:
@@
$\mathcal{L}_{\text{corr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\{y_{i}=\hat{y}_{i}\}\,\text{CE}(p_{i},y_{i})$ (2)
$\mathcal{L}_{\text{incorr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\{y_{i}\neq\hat{y}_{i}\}\,\text{KL}(p_{i}\parallel\mathcal{U})$ (3)
where $p_{i}$ is the predicted distribution for input $\mathbf{x}_{i}$, $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted label, $\mathcal{U}$ is the uniform distribution over all classes, and $N$ is the batch size. The cross-entropy term reduces the entropy of correct predictions, while the KL term increases entropy of incorrect ones.
The scalar $\alpha\in(0,1)$ balances the two terms. Smaller values emphasize penalizing overconfident errors (increasing deferral conservativeness), while larger values sharpen correct predictions and improve predictive accuracy. Thus, $\alpha$ directly controls the trade-off between deferral reliability and standalone model utility.
This idea is inspired by the OE loss for out-of-distribution detection (Hendrycks et al., 2018), which encourages uniform predictions on outlier data. However, to our knowledge, this formulation has not been used for improving deferral in model cascades.
Extension to Token-Based Models. For sequence models (e.g., LMs or VLMs), we compute the loss token-wise:
$\mathcal{L}_{\text{corr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathds{1}\{y_{i,t}=\hat{y}_{i,t}\}\,\text{CE}(p_{i,t},y_{i,t})$ (4)
$\mathcal{L}_{\text{incorr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathds{1}\{y_{i,t}\neq\hat{y}_{i,t}\}\,\text{KL}(p_{i,t}\parallel\mathcal{U})$ (5)
where $T$ is the sequence length, and $p_{i,t}$ is the token distribution at position $t$. This ensures fine-grained calibration across the sequence.
Stage 3: Confidence-Based Deferral. After fine-tuning with Gatekeeper, we use calibrated confidence scores from $\mathcal{M}_{S}$ to make deferral decisions. Following the selective prediction framework (El-Yaniv & Wiener, 2010), we define a gating function $g(\mathbf{x})$ and threshold $\tau$ for routing:
$(\mathcal{M}_{S},\mathcal{M}_{L},g)(\mathbf{x})=\begin{cases}\mathcal{M}_{S}(\mathbf{x})&\text{if }g(\mathbf{x})\geq\tau\\ \mathcal{M}_{L}(\mathbf{x})&\text{otherwise.}\end{cases}$ (6)
For classification models, we use max-softmax confidence:
$g_{\text{CL}}(\mathbf{x})=\max_{c}p(y=c\mid\mathbf{x})$ (7)
For sequence models, we use negative predictive entropy averaged across tokens:
$g_{\text{NENT}}(\mathbf{x})=\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{C}p(y_{t}=c\mid\mathbf{x})\log p(y_{t}=c\mid\mathbf{x})$ (8)
In both cases, higher scores indicate higher confidence. By thresholding this signal, we control which inputs $\mathcal{M}_{S}$ handles autonomously and which are escalated to $\mathcal{M}_{L}$—enabling an efficient, robust cascaded inference system.